OpenStudy (anonymous):

Need help!!! A party platter contains a viarety of dried fruit.To serve 16 people,96 pecies are needed how many fruit are needed to serve 22 people??

OpenStudy (anonymous):

16/96 = 22/x 16x = 22*96 x = 132